### v2

@@ 1 Introduction
The United Nations Educational, Scientific and Cultural Organization (UNESCO) aims to promote the identification, protection, and preservation of cultural and natural heritage worldwide, recognizing sites of outstanding value to humanity [^1] https://whc.unesco.org/en/conventiontext. Nowadays, the use of immersive technologies in cultural heritage displays has increasingly become the focus of academic research, as institutions seek to enhance visitor engagement and broaden access to valuable artifacts and experiences [1]. These immersive displays combine technologies such as virtual reality (VR), augmented reality (AR), mixed reality (MR), and other platforms like 360° video and interactive digital installations, enabling audiences to experience cultural heritage in more vivid, intuitive, and personal ways. However, the adoption of these technologies also introduces several risks and challenges.
@@
In the realm of immersive technologies for cultural heritage, prior studies have explored a variety of issues. Some research emphasizes digital preservation and conservation [2], while others focus on specific applications of immersive technologies or isolated aspects of cultural heritage [3, 4, 5, 6, 7, 8, 9, 10, 3, 11, 12]. However, certain works neglect to address the challenges or negative impacts—often referred to as the “dark sides”—associated with these technologies [13, 14, 15, 16, 17]. Additionally, in many cases, the application of immersive technologies to cultural heritage is restricted to VR, AR, and MR [18]. Furthermore, some studies rely on samples that predate 2018 [1, 19], or their sample sizes are too small to draw broader conclusions [19].
@@
Given these gaps, it is necessary to conduct a new survey of the literature on immersive technologies for cultural heritage, focusing on developments from the past five years. This review will place particular emphasis on the challenges and darker aspects of applying immersive technologies to digital cultural heritage, offering a critical reflection on their impact in this field.
@@
In this paper, we set the stage for a comprehensive review of the literature on immersive technologies in cultural heritage. Our review encompasses a pool of 5,368 articles published up to January 27, 2024, sourced from leading databases including the ACM Digital Library, IEEE Xplore, and Scopus. From this, we systematically analyzed 177 papers to identify these papers according to three main categories: “device,” “application,” and “technology.” Based on our analysis, we identified and summarized the strengths of the equipment, applications, and technologies, which can be grouped into the following areas: enhancing participation and experience, fostering social engagement, advancing education and knowledge dissemination, optimizing technological implementation, and promoting heritage protection and transmission.
@@
We also identified and categorized the shortcomings of the equipment, applications, and technologies, focusing on aspects such as technical barriers, sustainability, user experience, and safety concerns. Additionally, we found certain areas that have not been fully explored in the research, which could have a significant impact on the digital display and dissemination of cultural heritage. These gaps include the potential misinterpretation of cultural heritage, the influence on media purity, and the possible destructiveness to cultural heritage.
@@ 2.1 Research Status of Cultural Heritage Immersion Technology
Evens and Hauttekeete identify four major issues hindering the sustainability of digital preservation in cultural heritage institutions [2]. Boboc and Chong focus on the application of augmented reality to cultural heritage, with Boboc identifying eight key topics [3, 4], while Buragohain explores the challenges and strategies of digitization in the metaverse [5]. Idris discusses the technical challenges of protecting intangible cultural heritage [6], and Koller examines the research challenges in digital archives for 3D cultural relics [7]. Beraldin highlights the opportunities and challenges introduced by recent 3D technologies for heritage conservation and virtual reconstruction [8], while Windhager tackles the difficulties in visualizing collected data [9], and Baratin examines issues related to disseminating technical content [10].
@@ 2.2 Application of Immersive Technology in Cultural Heritage
The application of immersive technologies such as VR, AR, MR, and 360° video experiences has become an increasingly important area of research in cultural heritage, with a focus on enhancing visitor engagement and expanding access to artifacts and experiences. This section reviews key literature, covering the use of immersive technology in environmental reconstruction, data protection, visitor experience, education, and storytelling.
@@
On a larger scale, Argyriou et al. combine 360° immersive video with gamification and narrative-driven virtual experiences to create interactive urban environments. They argue that combining high participation through narrative and freedom through movement technology leads to deeper, long-lasting audience engagement [26].
@@ 2.3 Challenges and Limitations
Ethical concerns, particularly regarding cultural appropriation and the commercial use of digital heritage, have also arisen. Klinowski and Szafarowicz discuss conflicts between copyright, ownership, and public versus commercial interests in the digitization of cultural heritage [29]. Balancing the rights of creators and owners with public access and commercial use continues to be a significant challenge.
@@ 2.4 Research Gaps and Emergent Strategies
Despite the growing literature on the use of immersive technologies in various fields, there are still significant gaps in comprehensive, systematic reviews that focus specifically on their use in cultural heritage. Previous surveys have often touched on the potential and benefits of these technologies, but have largely ignored systematic exploration of the associated challenges and dark aspects, namely the unintended consequences of digitization on cultural heritage. These include, but are not limited to, issues of digital dependence and resistance; Social ethics in the digital context; The misinterpretation of cultural heritage may be brought about by the diverse digital display of cultural heritage; The impact of digital technology on the “purity of media experience” in the presentation of cultural heritage; The potentially damaging effects that digital display technologies can have on cultural heritage include physical damage, light and noise pollution, historical and cultural distortions, and threats to privacy and data security.
Existing reviews of the relationship between immersive technology, cultural heritage and people have often focused on specific aspects of research on how technology helps people to enhance their understanding of cultural heritage, but in our view this is not a comprehensive definition of the relationship between the three. For example, Sarah V. Bentley et al. describe the relationship between digital technology, the field of artificial intelligence, and people [30]. They identify social factors, as well as digital engagement factors under the influence of digital technologies and digital perception and experience factors as digital confidence that can influence people’s attitudes toward AI. Throughout the paper, the definition of the relationship between the three is expressed as a pattern of interdependence, which is often overlooked in the study of immersive technology and cultural heritage. Similarly, future research should focus on digital dependence and resistance under the influence of immersive technology, which may have an impact on the understanding of cultural heritage.
@@ 3 Method
With the aim of identifying and examining the latest status of research in the digital applications of cultural heritage up to 2024, we chose to conduct a scoping review. The purpose of the scope review is to “provide an initial delineation and assessment of the potential size and coverage of the existing research literature”. Although the issues concerned in systematic reviews are usually relatively precise, for example, the existing reviews in relevant fields usually focus on the current status of the application of cultural heritage in a specific field, which is a scope-defining review, lacking a comprehensive and systematic research perspective exploration that includes all the applications of cultural heritage technologies. They will clearly present and organize the relevant field evidence found, and analyze existing research and contributions under specific topics. But some authors argue that some kind of horizontal quality assessment should be done to better identify key gaps in the field. In contrast, we conduct research that is up-to-date in time and relatively larger in scope, rather than proposing solutions to specific problems. For this reason, we limit our literature screening to the full text of papers and scope review in the technologies, applications and devices of cultural heritage.
@@ 3.4 Filtering by Article Content
EC2. Key terms missing: If relevant key terms (e.g., augmented reality, virtual reality, immersive technologies, etc.) are not mentioned in the article, they are excluded. This reflects that the article may not be relevant to the research topic.
EC3. Terminology misuse: If the description or application of digital technology in the article is not consistent with the context in which cultural heritage is displayed, it is excluded. For example, if the “metadata” discussed in the article is used only for general information management or data storage, and does not relate to its application in the presentation of cultural heritage, such as to enhance the accessibility of cultural heritage materials, improve the interactivity of content or enrich the visitor experience, then such use is considered a misuse of the term. Because it does not properly reflect the important role and potential value of metadata in the display of cultural heritage. Similarly, if the technology is used for gamified learning or the simple presentation of cultural heritage content without an in-depth discussion of its specific application and impact in the presentation of cultural heritage, the article does not meet the research criteria.
EC4. Lack of correlation between digital technology and cultural heritage display: If the application of digital technology in the field of cultural heritage display is not explicitly discussed in the article, even if digital technology (e.g., augmented reality, virtual reality, artificial intelligence, etc.) is mentioned, it will be excluded. This is because a reference to digital technology without exploring its specific application and impact in the display of cultural heritage does not meet the core concerns of this study. For example, articles that focus on the protection of cultural heritage rather than its display, or discuss technologies that have a direct link to the display of cultural heritage, will also be excluded.
@@
EC6. Data set introduction: If the article mainly introduces data sets but does not explicitly mention augmented reality or artificial intelligence, the article will be excluded.
@@
IC1. Innovative applications of technology: Articles should show how emerging technologies (e.g., artificial intelligence, augmented reality, virtual reality, etc.) can be applied to the presentation and experience of cultural heritage. This includes, but is not limited to, innovative applications of technology in cultural heritage, ways to enhance interactive experiences, and the application of new technologies in cultural heritage education.
IC2. Integration of technology and cultural heritage: Research should explore new ways of integrating technology and cultural heritage, such as recreating historical scenes through augmented reality technology, or using virtual reality technology to display cultural sites that are not physically accessible.
@@
IC4. Application-oriented research: Consider those studies in which the application of technology is not a primary research focus but plays a key role in the presentation of cultural heritage. For example, a project that uses augmented reality to aid history education, or a study that applies machine learning techniques to analyze cultural heritage data.
@@ 4.1.1 Research Direction
Second, methodological framework research, mainly discusses the solution of a kind of cultural heritage digital display problems. For example, Suominen et al propose an approach to play culture’s internal and external cultural heritage, discussing the relationship between history, cultural heritage and digital technology through a four-fold table of the relationship between cultural heritage (or history) and digital technology [33]. Third, discuss the application of a new technology in the application of multiple projects. Ardito et al., for example, propose ways to identify and specify the attributes of IoT devices through user-defined semantics in order to make it easier for cross-domain experts to operate the technology and apply Internet of Things (IoT) technology more smoothly in the field of cultural heritage [34]. Fourth, the smallest proportion of literature review studies. Hou et al., for example, outline the models, techniques, and practices that drive the digitized life of intangible cultural heritage resources, looking at several key but less researched tasks such as digital archiving, computational coding, conceptual representation, and interactive engagement with intangible cultural elements to identify advancements and gaps in existing conventions  [35].
@@ 4.1.2 Publication Venue
A focus on publication venues can quickly determine the distribution of core journals within the field. According to the cumulative number of published papers, these papers were published in 5 publishing places, among which the most common ones were Journal on Computing and Cultural Heritage(35), Journal of Cultural Heritage(22), Journal of Cultural Heritage(35). It is preliminarily indicated that the above institutions are the main sources of contribution to the research in this field. The rest are relatively few: Digital Applications in Archaeology and Cultural Heritage(12), Personal and Ubiquitous Computing(12), and Heritage Science (11).
@@ 4.2.1 Sample Classification Analysis
We have carefully read and studied the full text of the 177 papers finally selected. We studied these papers according to three main categories: “device”, “application”, and “technology”.
Device refers to the devices used by the audience in the usage scenario mentioned in the paper. For example, Hulusic mentioned in the article that Tangible User Interfaces (TUI), in which users interact with the virtual cultural heritage application through the actual operation of the interface, can be summarized in the “mobile/PC/LAPTOP” category [36]. According to the data analysis, we mainly extracted a total of 199 frequency articles mentioned about equipment in the article. These include AR and VR devices, touchable and interactive user interface devices, mobile and portable devices, data acquisition scanning and printing devices, sensors and computing devices, projection and mapping technology devices, photography and video capture. Among them, the most widely used are AR and VR devices, which are mentioned 58 times, accounting for 29.1%. VR devices account for the largest proportion. In addition, data acquisition, 3D scanning and printing equipment were also mentioned 45 times, accounting for 22.6 percent. 3D laser scanners and photogrammetry equipment were used 22 times. In contrast, there are fewer articles that mention the use of touchable and interactive user interfaces, projection and mapping technology devices, photography and video capture devices, and we believe that there is still a lot of room for development in this field of devices (see Table 1).
@@
Application refers to the application field after the project development mentioned in the paper. For example, Tsiviltidou’s article explores the case of using digital storytelling in the classroom to construct inquiry-based learning with digital museum collections, a project applied to classroom instruction and thus grouped under the type of application “Education and learning” [95]. According to the data analysis, we mainly extracted a total of 195 applications mentioned in the article, including education and learning, community and social participation, digitization and visualization research, virtual and augmented reality technology applications, exhibition and museum experience enhancement, material and intangible cultural heritage protection and management, care for special groups, a total of 7 application fields. It is worth noting that 61 articles mentioned the field of exhibition and museum experience enhancement, accounting for 31.2%. There is a similar performance in the field of education and learning, accounting for 24.1 percent. It can be found that the application direction of the digital display of cultural heritage is still mainly to improve the basic function of user experience, while helping users better understand and learn the exhibition content, which is also the responsibility of museums. In addition, it is worth considering that there are relatively few applications in the field of community and social engagement, accounting for 9%. Therefore, we believe that in the process of digital display of cultural heritage, we should further open up the integration with the society and give play to the social functions of museums. In addition, the protection of cultural heritage and the care of special groups are the areas with the least attention, which also reflects the gap in the application of cultural heritage digitization in this field to a certain extent (see Table 2).
@@
4.2.2 Analysis of the Countries and Regions Involved
We counted the frequency of cultural heritage areas mentioned in 177 articles, and we found a relatively uneven distribution(see Figure 3. For example, China and Italy were mentioned more than 10 times as cultural heritage areas, with 21 and 15 mentions respectively. It was followed by the United Kingdom, the United States, Greece, Turkey and Poland, with four times each. We found that areas with more references to cultural heritage had relatively high levels of economic development and were relatively more open geographically, such as proximity to oceans or lakes. Spain was also mentioned four times as a cultural heritage site, while Jordan, Rome and Australia were each mentioned twice. These areas radiate around areas with a concentration of cultural heritage sites, and there may be relatively more cultural exchanges between these countries. Indonesia, Japan, Iraq, Russia and other countries are only mentioned once as cultural heritage sites, and these areas may be limited by relatively unstable environmental conditions, such as climatic environment, geological environment, international environment, etc., which is less likely to produce clustered cultural heritage (see Figure 3).
Figure 3: Analysis of the Country and Region Involved in the Literature Sample

### v3

@@ 1 Introduction
The United Nations Educational, Scientific and Cultural Organization (UNESCO) aims to promote the identification, protection, and preservation of cultural and natural heritage worldwide, recognizing sites of outstanding value to humanity [^1] https://whc.unesco.org/en/conventiontext. Nowadays, the use of immersive technologies in cultural heritage displays has increasingly become the focus of academic research, as institutions seek to enhance visitor engagement and broaden access to valuable artifacts and experiences [1]. These immersive displays combine technologies such as virtual reality (VR), augmented reality (AR), mixed reality (MR), and other platforms like 360-degree video and interactive digital installations, enabling audiences to experience cultural heritage —here encompassing both tangible and intangible forms— in more vivid, intuitive, and personal ways. However, the adoption of these technologies also introduces several risks and challenges.
@@
In the realm of immersive technologies for cultural heritage, prior studies have explored a variety of issues. Some research emphasizes digital preservation and conservation [2], while others focus on specific applications of immersive technologies or isolated aspects of cultural heritage [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, certain works neglect to address the challenges or negative impacts—often referred to as the “dark sides”—associated with these technologies [13, 14, 15, 16, 17]. Additionally, in many cases, the application of immersive technologies to cultural heritage is restricted to VR, AR, and MR [18]. Furthermore, some studies rely on samples that predate 2018 [1, 19], or their sample sizes are too small to draw broader conclusions [19].
@@
Given these gaps, it is necessary to conduct a new survey of the literature on immersive technologies for cultural heritage. This review will place particular emphasis on the grand challenges of applying immersive technologies to digital cultural heritage, offering a critical reflection on their impact in this field.
@@
In this paper, we set the stage for a comprehensive review of the literature on immersive technologies in cultural heritage. Our review covers a pool of 5,368 articles published up to January 27, 2024, sourced from leading databases including the ACM Digital Library, IEEE Xplore, and Scopus. From this, we systematically analyzed 177 papers to identify these papers according to three main categories: “device,” “application,” and “technology.” Based on our analysis, we identified and summarized the strengths of the equipment, applications, and technologies, which can be grouped into the following areas: enhancing participation and experience, fostering social engagement, advancing education and knowledge dissemination, optimizing technological implementation, and promoting heritage protection and transmission.
@@
We also identified and categorized the shortcomings of the equipment, applications, and technologies, focusing on aspects such as technical barriers, sustainability, user experience, and safety concerns. Additionally, we found certain areas that have not been fully explored in the research, which could have a significant impact on the digital display and dissemination of cultural heritage. These gaps include the potential misinterpretation of cultural heritage, the influence on media purity, and the possible potential damage to cultural heritage.
@@ 2.1 Research Status of Cultural Heritage Immersion Technology
Evens and Hauttekeete identify four major issues hindering the sustainability of digital preservation in cultural heritage institutions [2]. Boboc and Chong focus on the application of AR to cultural heritage, with Boboc identifying eight key topics [3, 4], while Buragohain explores the challenges and strategies of digitization in the metaverse [5]. Idris discusses the technical challenges of protecting intangible cultural heritage [6], and Koller examines the research challenges in digital archives for 3D cultural relics [7]. Beraldin highlights the opportunities and challenges introduced by recent 3D technologies for heritage conservation and virtual reconstruction [8], while Windhager tackles the difficulties in visualizing collected data [9], and Baratin examines issues related to disseminating technical content [10].
@@ 2.2 Application of Immersive Technology in Cultural Heritage
The application of immersive technologies such as VR, AR, MR, and 360-degree video experiences has become an increasingly important area of research in cultural heritage, with a focus on enhancing visitor engagement and expanding access to artifacts and experiences. This section reviews key literature, covering the use of immersive technology in environmental reconstruction, data protection, visitor experience, education, and storytelling.
@@
On a larger scale, Argyriou et al. combine 360-degree immersive video with gamification and narrative-driven virtual experiences to create interactive urban environments. They argue that combining high participation through narrative and freedom through movement technology leads to deeper, long-lasting audience engagement [26].
@@ 2.3 Challenges and Limitations
Regarding cultural appropriation and the commercial use of digital heritage have also arisen. Klinowski and Szafarowicz discuss conflicts between copyright, ownership, and public versus commercial interests in the digitization of cultural heritage [29]. This illustrates that balancing the rights of creators and owners with public access and commercial use could still be a major challenge. However, the potential of digital approaches to promote wider access and create business opportunities still requires validation through further empirical research.
@@ 2.4 Research Gaps and Emergent Strategies
Despite the growing literature on the use of immersive technologies in various fields, there are still significant gaps in comprehensive, systematic reviews that focus specifically on their use in cultural heritage. Previous surveys have often touched on the potential and benefits of these technologies, but have largely overlooked a systematic exploration of the associated challenges and negative aspects, particularly the unintended consequences of digitization on cultural heritage. These include, but are not limited to, issues such as digital dependence and resistance, social ethics in the digital context, and the misinterpretation of cultural heritage caused by its diverse digital representations. Another concern is the impact of digital technology on the “purity of media experience” in presenting cultural heritage. This concept refers to the authenticity and immediacy with which audiences engage with cultural heritage. Ideally, this experience should reflect the original historical and sensory context, without significant mediation or alteration. When individuals engage with cultural heritage, the goal is for them to encounter it in its unaltered form, minimally influenced by interpretive layers, technological enhancements, or information filtering. This principle emphasizes that the processes of media translation and interpretation should not substantially distort the original essence of cultural heritage. Further, the use of digital display technologies in cultural heritage can have potentially damaging effects, including physical damage, light and noise pollution, historical and cultural distortions, and threats to privacy and data security.
Existing reviews of the relationship between immersive technology, cultural heritage and people have often focused on specific aspects of research on how technology helps people to enhance their understanding of cultural heritage, but in our view this is not a comprehensive definition of the relationship between the three. For example, Bentley et al. describe the relationship between digital technology, the field of artificial intelligence, and people [30]. They identify social factors, as well as digital engagement factors under the influence of digital technologies and digital perception and experience factors as digital confidence that can influence people’s attitudes toward AI. Throughout the paper, the definition of the relationship between the three is expressed as a pattern of interdependence, which is often overlooked in the study of immersive technology and cultural heritage. Similarly, future research should focus on digital dependence and resistance under the influence of immersive technology, which may have an impact on the understanding of cultural heritage.
@@ 3 Method
With the aim of identifying and examining the latest status of research in the digital applications of cultural heritage up to 2024, we chose to conduct a scoping review. The purpose of the scope review is to “provide an initial delineation and assessment of the potential size and coverage of the existing research literature”. Although the issues concerned in systematic reviews are usually relatively precise, for example, the existing reviews in relevant fields usually focus on the current status of the application of cultural heritage in a specific field, which is a scope-defining review, lacking a comprehensive and systematic research perspective exploration that includes all the applications of cultural heritage technologies. They will clearly present and organize the relevant field evidence found, and analyze existing research and contributions under specific topics. In this paper, we conduct research that is up-to-date in time and relatively larger in scope, rather than proposing solutions to specific problems. For this reason, we limit our literature screening to the full text of papers and scope review of the technologies, applications and devices of cultural heritage.
@@ 3.4 Filtering by Article Content
EC2. Key terms missing: If relevant key terms (e.g., AR, VR, immersive technologies, etc.) are not mentioned in the article, they are excluded. This reflects that the article may not be relevant to the research topic.
EC3. Inappropriate Use of Terminology: If the description or application of digital technology in the article is not consistent with the context in which cultural heritage is displayed, it is excluded. For example, if the “metadata” discussed in the article is used only for general information management or data storage, and does not relate to its application in the presentation of cultural heritage, such as to enhance the accessibility of cultural heritage materials, improve the interactivity of content or enrich the visitor experience, then such use is considered an inappropriate use of the term. Because it does not properly reflect the important role and potential value of metadata in the display of cultural heritage. Similarly, if the technology is used for gamified learning or the simple presentation of cultural heritage content without an in-depth discussion of its specific application and impact in the presentation of cultural heritage, the article does not meet the research criteria.
EC4. Lack of correlation between digital technology and cultural heritage display: If the application of digital technology in the field of cultural heritage display is not explicitly discussed in the article, even if digital technology (e.g., AR, VR, etc.) is mentioned, it will be excluded. This is because a reference to digital technology without exploring its specific application and impact in the display of cultural heritage does not meet the core concerns of this study. For example, articles that focus on the protection of cultural heritage rather than its display, or discuss technologies that have a direct link to the display of cultural heritage, will also be excluded.
@@
EC6. Data set introduction: If the article mainly introduces data sets but does not explicitly mention immersive or emerging technologies, the article will be excluded.
@@
IC1. Innovative applications of technology: Articles should show how emerging technologies (e.g., artificial intelligence, AR, VR, etc.) can be applied to the presentation and experience of cultural heritage. This includes, but is not limited to, innovative applications of technology in cultural heritage, ways to enhance interactive experiences, and the application of new technologies in cultural heritage education.
IC2. Integration of technology and cultural heritage: Research should explore new ways of integrating technology and cultural heritage, such as recreating historical scenes through AR, or using VR to display cultural sites that are not physically accessible.
@@
IC4. Application-oriented research: Consider those studies in which the application of technology is not a primary research focus but plays a key role in the presentation of cultural heritage. For example, a project that uses AR to aid history education, or a study that applies machine learning techniques to analyze cultural heritage data.
@@
For the process in the literature screening, we employed a multi-step approach that included independent screening by multiple authors, as well as a review by a panel of 20 experts. These experts, comprising cultural heritage presentation practitioners, university scholars, and professionals from cultural heritage protection institutions, were invited to assess and discuss the initial selection results. Their involvement enhances both the objectivity and credibility of the screening process. Additionally, an anonymized table (see Table 1) will be included in the manuscript, providing background information on these 20 experts, such as their role, gender, years of experience (categorized as senior [>15absent15>15> 15 years] or junior [<5absent5<5< 5 years]), and professional background (i.e., academic, industry, and public sector). This information is included to ensure greater transparency regarding the composition and expertise of the panel.
Table 1: Anonymized Background Information of the 20 Experts in the Review Panel
Expert Gender Age Role Background Experience
1 Male 55 Professor Academic Senior
2 Female 36 PhD Student Academic Senior
3 Female 31 Historian Academic Junior
4 Male 33 Architectural Restorer Academic Junior
5 Male 46 Cultural Creative Company Owner Industry Senior
6 Female 41 Heritage Site Manager Public Sector Senior
7 Male 34 Heritage Site Interpreter Public Sector Senior
8 Male 29 Local Museum Staff Public Sector Junior
9 Male 30 National Museum Staff Public Sector Junior
10 Female 35 Cultural Heritage Column Website Manager Industry Junior
11 Female 36 Cultural Heritage Column Reporter Industry Senior
12 Male 37 Associate Professor Academic Senior
13 Male 38 Local Museum Staff Public Sector Senior
14 Male 44 Vice President of Cultural Creative Group Industry Senior
15 Male 45 Local Museum Staff Public Sector Senior
16 Female 40 Local Museum Staff Public Sector Senior
17 Male 44 Local Museum Staff Public Sector Senior
18 Female 35 Local Museum Staff Public Sector Junior
19 Male 40 Researcher Academic Senior
20 Female 33 Local Museum Staff Public Sector Junior
@@ 4.1.1 Research Direction
Second, methodological framework research, mainly discusses the solution of a kind of cultural heritage digital display problems. For example, Suominen et al propose an approach to play culture’s internal and external cultural heritage, discussing the relationship between history, cultural heritage and digital technology through a four-fold table of the relationship between cultural heritage (or history) and digital technology [33]. Third, discuss the application of a new technology in the application of multiple projects. Ardito et al., for example, propose ways to identify and specify the attributes of Internet of Things (IoT) devices through user-defined semantics in order to make it easier for cross-domain experts to operate the technology and apply IoT technology more smoothly in the field of cultural heritage [34]. Fourth, the smallest proportion of literature review studies. Hou et al., for example, outline the models, techniques, and practices that drive the digitized life of intangible cultural heritage resources, looking at several key but less researched tasks such as digital archiving, computational coding, conceptual representation, and interactive engagement with intangible cultural elements to identify advancements and gaps in existing conventions [35].
@@ 4.1.2 Publication Venue
A focus on publication venues can quickly determine the distribution of core journals within the field. According to the cumulative number of published papers, these papers were published in several publishing places, among which the most common five were Journal on Computing and Cultural Heritage (35), Journal of Cultural Heritage (22), Digital Applications in Archaeology and Cultural Heritage (12), Personal and Ubiquitous Computing (12), and Heritage Science (11).
@@ 4.2.1 Sample Classification Analysis
We have carefully read and studied the full text of the 177 papers selected finally. We studied these papers according to three main categories: “device”, “application”, and “technology”.
Device refers to the devices used by the audience in the usage scenario mentioned in the paper. For example, Hulusic mentioned in the article that Tangible User Interfaces (TUI), in which users interact with the virtual cultural heritage application through the actual operation of the interface, can be summarized in the “mobile/PC/LAPTOP” category [36]. According to the data analysis, we mainly extracted a total of 199 frequency articles mentioned about equipment in the article. These include AR and VR devices, touchable and interactive user interface devices, mobile and portable devices, data acquisition scanning and printing devices, sensors and computing devices, projection and mapping technology devices, photography, and video capture. Among them, the most widely used are AR and VR devices, which are mentioned 58 times, accounting for 29.1%. VR devices account for the largest proportion. In addition, data acquisition, 3D scanning and printing equipment were also mentioned 45 times, accounting for 22.6 percent. 3D laser scanners and photogrammetry equipment were used 22 times. In contrast, there are fewer articles that mention the use of touchable and interactive user interfaces, projection and mapping technology devices, photography and video capture devices, and we believe that there is still a lot of room for development in this field of devices (see Table 2).
@@
Application refers to the application field after the project development mentioned in the paper. For example, Tsiviltidou’s article explores the case of using digital storytelling in the classroom to construct inquiry-based learning with digital museum collections, a project applied to classroom instruction and thus grouped under the type of application “Education and learning” [95]. According to the data analysis, we mainly extracted a total of 195 applications mentioned in the article, including education and learning, community and social participation, digitization and visualization research, virtual and AR technology applications, exhibition and museum experience enhancement, material and intangible cultural heritage protection and management, care for special groups, a total of 7 application fields. It is worth noting that 61 articles mentioned the field of exhibition and museum experience enhancement, accounting for 31.2%. There is a similar performance in the field of education and learning, accounting for 24.1 percent. It can be found that the application direction of the digital display of cultural heritage is still mainly to improve the basic function of user experience, while helping users better understand and learn the exhibition content, which is also the responsibility of museums. In addition, it is worth considering that there are relatively few applications in the field of community and social engagement, accounting for 9%. Therefore, we believe that in the process of digital display of cultural heritage, we should further open up the integration with the society and give play to the social functions of museums. In addition, the protection of cultural heritage and the care of special groups are the areas with the least attention, which also reflects the gap in the application of cultural heritage digitization in this field to a certain extent (see Table 3).